# Pretrained Hybrids with MAD Skills

### v1

@@ 3.2 How To Use Manticore
With Manticore, we can automatically select language models without training every model in the search space, automatically construct pretrained hybrid architectures without significant trial-and-error, and also program pretrained hybrids without full training. In this section, we will discuss the details of how Manticore can be used in each of these three usage scenarios.
@@
Training hybrids from scratch. Manticore can be used to automatically select LMs without training all of the LMs in the search space. Our selection technique is simple: inspired by gradient-based NAS techniques [23] and treating the mixture weights as our ‘architecture parameters,’ we proceed in two steps: 1. train mixture weights along with all other parameters, and 2. freeze the mixture weights and retrain the rest of the parameters from scratch. Unlike NAS, we found that in many pretraining settings, it was sufficient to stop at 1. and forgo retraining. In our pretraining experiments, we primarily use randomly-initialized GPT-Neo [5] and Mamba [13] as component models without projectors, and separately experiment with a subset of the blocks from MAD [34].
@@
Fine-tuning pretrained hybrids. Manticore can be used to create and fine-tune pretrained hybrids. We create pretrained hybrids as follows: begin with a set of pretrained models, replace their LM heads and embeddings with a single randomly initialized LM head and embedding layer, and pretrain the projectors on a small amount of general language data such as FineWeb [30] while keeping the original component model weights frozen.444 We found one billion tokens to be sufficient for projector pretraining. To fine-tune the pretrained hybrids on downstream task data, we first search for mixture weights by training all of the parameters simultaneously, we freeze the mixture weights, rewind the component models and projectors to their pretrained state, and fine-tune. This procedure completely sidesteps large-scale pretraining of new hybrids. In our synthetic experiments, we create pretrained Manticore hybrids from pretrained GPT-Neo-125M [5] and Mamba-130M [13] models, while for our experiments on real natural language data, we opt for pretrained Pythia-410M [4] and Mamba-370M [13] as component models.
@@
Programming hybrids. Excitingly, we can program Manticore mixture weights by using external information to predict them. We consider two scenarios. If we know that a component model has blocks that are somehow incompatible with the target task—e.g. resulting from sequence length constraints—we can omit these blocks by setting their mixture weights to 0. Otherwise, we can predict good mixture weights by searching on a fixed set of proxy tasks—for this, we use the MAD tasks [34]. The MAD tasks are synthetic unit tests that are predictive of hybrid LM scaling laws, but within our framework, we find that they are also useful for finding general-purpose pretrained hybrids. We use the following procedure for programming mixture weights using the MAD tasks. First, run search on the MAD tasks using a smaller, randomly initialized version of our pretrained hybrid. For each MAD task, our search procedure returns a set of mixture weights—we simply average the resulting mixture weights, freeze them, and fine-tune on the downstream task data.
@@ 4 Experimental Results
  • C2. Trained from scratch, Manticore hybrids are competitive with existing hybrids and LMs, and
  • C3. We can program mixture weights using external sources without search on the task data.
@@ 4.1 Fine-Tuning Pretrained Hybrids
We evaluate C1, first on a synthetic fine-tuning task, and then on natural language fine-tuning tasks.
Setup. We consider a synthetic LM dataset comprising GPT-Neo and Mamba generated completions of text from Penn Treebank [27]. Naturally, we also use pretrained GPT-Neo-125M and Mamba-130M models as component models, creating a single Manticore block with projectors that were pretrained on one billion tokens from FineWeb [30]. We perform search using DARTS and perform post-search retraining with the model weights and projectors rewound to their pretrained state.
Results. Our results are shown in Figure 5 (left). We compare our search results to a sweep over a range of possible mixture weights, and find that our search procedure returns the optimal mixture weights, outperforming both Mamba and GPT-Neo. This confirms our claim that Manticore hybrids can outperform their component models on synthetic fine-tuning tasks. Given that this task comprises two slices that each of our component models should be good at—GPT-Neo should be good at predicting GPT-Neo outputs, and vice versa—we hypothesize that Manticore hybrids are well suited to situations in which the component models have complementary ‘skills’ [7].
Figure 5: Mixture weight sweeps on Penn Treebank completions using pretrained GPT-Neo-125M and Mamba-130M as our component models. (Left) When we create one Manticore block, there is a region of the search space where we improve over Mamba. (Right) The same holds for two Manticore blocks, and our technique for hybrid programming using MAD discovers this region.
Setup. We evaluate on three natural language fine-tuning datasets: Penn Treebank [27], the Alpaca instructions dataset [37], and ELI5 [10]. We use Pythia-410M and Mamba-370M as component models, and create a single Manticore block from their blocks with projectors that were pretrained on one billion tokens from FineWeb [30]. As before, we first search for mixture weights, and then we retrain with the fixed mixture weights found by search.
Results. Our results are shown in Table 1. Manticore outperforms its component models on Alpaca and ELI5, while it achieves performance between its two component models on Penn Treebank. This confirms our claim that Manticore can outperform component models on real natural language tasks. The fact that Mamba-370M outperforms Manticore in this setting is not a failure of our framework, as Mamba-370M is included as part of our search space. We speculate that the use of more powerful search procedures from the NAS literature, such as GAEA [21], could improve our search performance and help to recover or outperform Mamba-370M.
Table 1: Manticore on natural language tasks using Pythia-410m and Mamba-370m as component models. The best test losses are bolded and the second-best are underlined.
@@ 4.2 Training Hybrids from Scratch
For C2, we compare to prior hybrids on MAD and non-hybrid component models on LRA and MAD.
Setup. We compare training Manticore from scratch to training existing hybrid architectures on the MAD tasks. We begin with two hybrid architectures from the literature: Mambaformer [29], which combines Mamba and attention blocks, and the striped multi-head Hyena + Mixture-of-Experts (MoE) MLP architecture that was shown to perform well on the MAD tasks [34]. We compare these two baselines to a Manticore hybrid combining three component models: striped multi-head Hyena + MoE-MLP, a transformer, and Mamba. We use two blocks for each of these architectures, creating two Manticore blocks. Again, we search for mixture weights and then retrain.
Results. The results of this experiment are shown in Table 2. We outperform the striped multi-head Hyena + MoE model from the MAD paper, and we approach the performance of Mambaformer on all but one task. This validates the claim that Manticore hybrids, trained from scratch, compete with existing hybrids. Despite Mambaformer not being a component model, it is in our search space, and we again speculate that improvements in search would lead to its recovery.
Task Striped MH Hyena Mambaformer Manticore
+ MoE-MLP
In-context Recall 3.7153 0.0020 0.0048
Fuzzy In-context Recall 4.1714 4.1712 4.1750
Noisy In-context Recall 4.1643 4.1646 4.1607
Selective Copying 1.8021 0.0005 0.0171
Memorization 8.8353 5.2179 8.9254
Setup. We compare Manticore hybrids to their component models on LRA, when trained from scratch. We create GPT-Neo and Mamba component models of similar sizes to those in Tay et al. [38] and create a Manticore hybrid. As a simplified pipeline, we do not retrain model weights after search.
Results. Our results are shown in Table 3. We outperform the component models on all tasks except for IMDb, in which case Manticore was between GPT-Neo and Mamba. This validates the claim that Manticore hybrids, trained from scratch, compete with existing LMs.
Table 3: Manticore hybrids trained from scratch on LRA using GPT-Neo and Mamba component models. The best test accuracies are bolded. GPT-Neo does not support the Pathfinder-X sequence length requirement, so we set its mixture weight to 0 and Manticore reduces to Mamba.
@@
Setup. Next, we compare Manticore to non-hybrid architectures trained from scratch on the MAD tasks. We compare two-block GPT-Neo and Mamba models to a Manticore hybrid using a single Manticore block. Again, we report the performance of the search procedure itself without retraining.
Results. Our results are shown in Table 4. Manticore outperforms GPT-Neo and Mamba on all of the MAD tasks in this setting. This provides further evidence for our claim that Manticore hybrids compete with existing LMs when trained from scratch. It is conceivable that our larger Manticore hybrids simply perform better than component models due to their size—however, we find that post-search discretization and retraining tends to result in similar performance, but reduces the model size by roughly half. We include an ablation of post-search discretization in the Appendix.
Table 4: Trained from scratch on the MAD tasks, Manticore improves over small two-block component models combined into a single Manticore block. The best test losses are shown in bold.
Task GPT-Neo (A) Mamba (B) Manticore [A, B]
In-context Recall 4.0771 4.1858 4.0768
Fuzzy In-context Recall 4.4384 4.8097 4.2797
Noisy In-context Recall 4.1843 4.2605 4.1823
Selective Copying 1.0470 3.7765 0.9478
Memorization 4.6110 5.2281 4.1367
@@ 4.3 Programming Hybrids
We evaluate C3 with two types of external data: access to task metadata such as sequence length requirements, and the use of the MAD tasks as a proxy for search on downstream task data.
@@
Setup. As in many of our previous experiments, we used the GPT-Neo and Mamba architectures as component models to our Manticore hybrid. However, this time, we set out to train from scratch on the extremely long-range Pathfinder-X task from LRA, which requires sequence length support greater than that of GPT-Neo. Using this external information about the task, we set the mixture weights for GPT-Neo to 0, which in this case, means that Manticore reduces to Mamba. [^5] As of writing, Mamba on LRA is open: https://github.com/state-spaces/mamba/issues/282.
@@
Results. The results of this experiment are shown in the last row of Table 3. In the simple case of having access to task metadata, this validates the claim that we can program mixture weights to exclude incompatible blocks. At the time of writing, we are not aware of prior published Mamba results on LRA despite community interest, which would make our evaluation in Table 3 the first such result. Note that we did not thoroughly tune hyperparameters, so we view this result as a starting point for the community, rather than a final answer.
@@
Setup. Finally, in the case in which we can run all of our component models on our learning task, we program the mixture weights using the MAD tasks as a search proxy. We set out to fine-tune a pretrained hybrid comprising GPT-Neo-125M and Mamba-130M with two Manticore blocks on our Penn Treebank completions synthetic. We train a scaled-down version of this Manticore hybrid with randomly initialized weights and two blocks per component model on the MAD tasks. This yields mixture weights for each of the MAD tasks—we average them and then fine-tune our pretrained hybrid on Penn Treebank completions using the predicted mixture weights.
@@
Results. Our results are shown in Figure 5 (right). We superimpose the predicted mixture weights and mean search trajectory from MAD onto the architecture loss landscape computed on Penn Treebank completions. We find that this procedure recovers a hybrid that outperforms the component models (Mamba, lower right; GPT-Neo, upper left) and substantially outperforms the naive frankenmerges in our search space (upper right and lower left) [12]. This validates the claim that we can program mixture weights using external sources without performing search on the task data. Intriguingly, search on the MAD tasks appears to follow the architecture gradient on the different downstream fine-tuning task, even though the architecture is scaled-down and trained from scratch on MAD. We suspect that the mixture weights and architecture loss landscapes for pretrained hybrids are fairly universal across fine-tuning tasks, and that the same procedure is likely to work on a wide range of scenarios. Furthermore, we hypothesize that this technique could outperform other gradient-based NAS methods directly applied to the downstream task.

### v2

@@ 3.2 How To Use Manticore
With Manticore, we can automatically select language models without training every model in the search space, automatically construct pretrained hybrid architectures without significant trial-and-error, and program pretrained hybrids without full training. In this section, we discuss the details of how Manticore can be used in each of these three usage scenarios.
@@
Training hybrids from scratch. Manticore can be used to automatically select LMs without training all of the LMs in the search space. Our selection technique is simple: inspired by gradient-based NAS techniques [26] and treating the mixture weights as our ‘architecture parameters,’ we proceed in two steps: 1. train mixture weights along with all other parameters, and 2. freeze the mixture weights and retrain the rest of the parameters from scratch. Unlike NAS, we found that in many pretraining settings, it was sufficient to stop at 1. and forgo retraining. In our pretraining experiments, we use randomly-initialized GPT-Neo [5] and Mamba [14] as component models without projectors, and separately experiment with a subset of blocks from MAD [37].
@@
Fine-tuning pretrained hybrids. Manticore can be used to create and fine-tune pretrained hybrids. We create pretrained hybrids as follows: begin with a set of pretrained models, replace their LM heads and embeddings with a single randomly initialized LM head and embedding layer, and pretrain the projectors on a small amount of general language data such as FineWeb [33] while keeping the original component model weights frozen. [^5] We found that 100M tokens sufficed for projector pretraining. To fine-tune the pretrained hybrids on downstream task data, we first search for mixture weights by training all of the parameters simultaneously, we freeze the mixture weights, rewind the component models and projectors to their pretrained state, and fine-tune. This procedure completely sidesteps large-scale pretraining of new hybrids. [^6] We include an extensive FLOPs analysis and a discussion of comparable baselines in the Appendix. In our synthetic experiments, we create pretrained Manticore hybrids from pretrained GPT-Neo-125M [5] and Mamba-130M [14] models, while for our experiments on real natural language data, we opt for pretrained Pythia-410M [4] and Mamba-370M [14] as component models.
@@
Programming hybrids. Excitingly, there are cases in which we can program Manticore mixture weights by using external information to predict them. We consider two scenarios. If we know that a component model has blocks that are incompatible with the target task—e.g. resulting from sequence length constraints—we can omit these blocks by setting their mixture weights to 0. Otherwise, we can predict good mixture weights by searching on a fixed set of proxy tasks. For this, we use MAD tasks [37]. The MAD tasks are synthetic unit tests that are predictive of hybrid LM scaling laws, but within our framework, we find that MAD can also be useful for finding pretrained hybrids. We use the following procedure for programming mixture weights using the MAD tasks. First, run search on the MAD tasks using a smaller, randomly initialized version of our pretrained hybrid. For each MAD task, our search procedure returns a set of mixture weights—we simply average the resulting mixture weights, freeze them, and fine-tune on downstream task data.
3.3 Discussion and Design Considerations
Manticore features several intentional design decisions that we make concrete in this section.
Where Manticore hybrids excel.
It is known that hybrids excel at compositional tasks like finding a token arbitrarily far in the past and then performing a local copy operation—this for instance, necessitates a tradeoff in SSM [15] state size and transformer context. Results like these motivate the study of tools like Manticore. For this reason, we expect Manticore to excel at tasks in which component models are specialized for certain data sources or aspects of the dataset. As a result, many of our experiments in Section 4 feature heterogeneous data sources.
Design tradeoffs in Manticore.
Manticore requires taking a forward pass with each of its component models, which increases inference cost over the use of a single component model. This increased inference cost is an explicit tradeoff for not having to pretrain a hybrid from scratch. In Appendix D, we motivate this tradeoff by showing that the total FLOPs required to produce a Manticore hybrid is dominated by component model pretraining, and that this can be avoided by reusing existing pretrained models. Due to the simplicity of our projector architecture, we also show that the inference cost of Manticore is dominated by forward passes of its component models. This further motivates our comparison to ensembles in Section 4, due to their similar inference and training FLOPs requirements. Finally, Manticore can be scaled to larger component models without significant overhead, as the inference costs scale linearly in the size of its component models.
Flexibility of search algorithm.
Our search space works best with NAS algorithms that support continuous-valued mixture weights, such as DARTS [26], GAEA [24], and other gradient-based NAS algorithms. This makes our framework particularly flexible in its support for this broad class of NAS algorithms, while leaving room for specialized algorithms to be developed later. In Appendix A, we include an ablation comparing DARTS to the DASH [39] search algorithm, along with various other components of the NAS pipeline. These ablations help characterize the desirable traits of NAS search algorithms for Manticore. For the purposes of our experiments, we mainly rely on DARTS [26]—an entirely off-the-shelf NAS algorithm—and leave the development of tailor-made hybrid search algorithms to future work.
@@ 4 Experimental Results
  • C2. Trained from scratch, Manticore is competitive with existing hybrids and LMs, and
  • C3. In certain cases, we can program mixture weights without search on the task data.
@@ 4.1 Fine-Tuning Pretrained Hybrids
We evaluate C1, first on a synthetic task, and then on natural language fine-tuning tasks.
Setup. We consider a synthetic LM dataset comprising GPT-Neo and Mamba generated completions of text from Penn Treebank [30]. Naturally, we also use pretrained GPT-Neo-125M and Mamba-130M component models, creating a single Manticore block with projectors that were pretrained on 100M tokens from FineWeb [33]. We search using DARTS, and afterward, rewind the model weights and projectors to their pretrained states for retraining.
Results. Our results are shown in Figure 2 (left). We compare our search results to a sweep over a range of possible mixture weights and find that our search procedure returns the optimal mixture weights, outperforming both Mamba and GPT-Neo. This confirms our claim that Manticore hybrids can outperform their component models on synthetic fine-tuning tasks. Given that this task comprises two slices that each of our component models should be good at—GPT-Neo should be good at predicting GPT-Neo outputs, and vice versa—we hypothesize that Manticore hybrids are especially well suited to the component models having complementary ‘skills’ [7].
Figure 2: Mixture weight sweeps on Penn Treebank completions using pretrained GPT-Neo-125M and Mamba-130M as our component models. (Left) When we create one Manticore block, there is a region of the search space where we improve over Mamba. Here, we denote the loss value and mixture weights found via search using a yellow star and track the loss throughout training in green. (Right) The same holds for two Manticore blocks, and our technique for hybrid programming using MAD discovers this region.
Setup. We evaluate on three natural language fine-tuning datasets: Penn Treebank [30], the Alpaca instructions dataset [40], and ELI5 [10]. We use Pythia-410M and Mamba-370M as our component models, and create a single Manticore block from the blocks of the two models with projectors that were pretrained on 100M tokens from FineWeb [33]. As before, we search for mixture weights and retrain with the fixed mixture weights found by search.
Results. Our results are shown in Table 1. Manticore outperforms its component models on Alpaca and ELI5, while it achieves performance between its two component models on Penn Treebank. This confirms our claim that Manticore can outperform component models on real natural language tasks. The fact that Mamba-370M outperforms Manticore in this setting is not a failure of our framework, as Mamba-370M is included as part of our search space—improving the search procedure beyond off-the-shelf NAS algorithms in order to obtain these high performing models is an interesting direction for future work.
@@
Table 1: Manticore on language tasks using Pythia-410m and Mamba-370m component models. The best test losses are bolded and the second-best are underlined.
Setup. Building on the previous setup for natural language tasks, we perform a sweep over the $\alpha$ parameter corresponding to Mamba in our search space, and compare the results of the sweep to off-the-shelf NAS algorithms: DARTS [26] (Manticore’s search procedure), GAEA [24], and DASH [39]. In order to compare Manticore to a method with comparable inference cost, we also consider an ensemble baseline where the ensemble weights are learned during training. For three datasets, 50% of the documents are drawn from the Alpaca [40] dataset to artificially induce heterogeneity–we hypothesize that Manticore hybrids are well-suited to such settings—if Manticore’s component models specialize in different subsets of a dataset, then Manticore should achieve improved overall performance on the combined dataset.
Results. Our results are shown in Figure 3. We find that in all but one setting (NI Chinese QA + Alpaca), at least two of the NAS algorithms that we evaluate recover a model that outperforms its component models. Furthermore, on five of the datasets, at least one NAS algorithm outperforms or matches the best model found during the sweep. Manticore also substantially outperforms the ensemble on all tasks. This is further evidence for our claim that Manticore outperforms component models on natural language, and demonstrates that NAS algorithms can find performant pretrained hybrids in our search space.
Figure 3: Mixture weight sweeps using Pythia-410M and Mamba-370M component models. NAS algorithms often locate regions of the search space that outperform component models and a learned ensemble baseline.
@@ 4.2 Training Hybrids from Scratch
For C2, we compare to prior hybrids on MAD and non-hybrid models on LRA and MAD.
Setup. We compare training Manticore from scratch to training existing hybrid architectures on MAD tasks. We begin with two hybrid architectures from the literature: Mambaformer [32], which combines Mamba and attention blocks, and the striped multi-head Hyena + Mixture-of-Experts (MoE) MLP architecture that was shown to perform well on the MAD tasks [37]. We compare these two baselines to a Manticore hybrid combining three component models: striped multi-head Hyena + MoE-MLP, a transformer, and Mamba. We use two blocks for each of these architectures, creating two Manticore blocks. Again, we search for mixture weights and then retrain.
Results. The results of this experiment are shown in Table 2 (left). We outperform the striped multi-head Hyena + MoE model from the MAD paper, and we approach the performance of Mambaformer on all but one task. This validates the claim that Manticore hybrids, trained from scratch, compete with existing hybrids. Despite Mambaformer not being a component model, it is in our search space, and we again speculate that improvements in search would lead to its recovery.
Task Starting from existing hybrids Starting from non-hybrids
SMH Hyena Mamba- Manticore GPT-Neo Mamba Manticore
+ MoE-MLP (A) former (B) (C) (D) [C, D]
Ctx. Recall 3.7153 0.0020 0.0048 4.0771 4.1858 4.0768
Fuzzy Recall 4.1714 4.1712 4.1750 4.4384 4.8097 4.2797
Noisy Recall 4.1643 4.1646 4.1607 4.1843 4.2605 4.1823
Sel. Copy 1.8021 0.0005 0.0171 1.0470 3.7765 0.9478
Mem. 8.8353 5.2179 8.9254 4.6110 5.2281 4.1367
Table 2: Results for training from scratch on MAD tasks. (Left) Manticore matches the performance of existing hybrids on all but one task. (Right) Manticore improves over non-hybrid component models. (Both) best losses are bolded and second best are underlined.
Setup. We compare Manticore hybrids to their component models on LRA, when trained from scratch. We use GPT-Neo and Mamba component models of similar sizes to those in Tay et al. [41] to create Manticore hybrids, while keeping the number of blocks the same between the component models. In these experiments, we create a Manticore block for every block in the component models, ranging from 3 to 6 Manticore blocks.
Results. Our results are shown in Table 3. We outperform component models on all tasks except for IMDb. This validates the claim that Manticore hybrids, trained from scratch, compete with existing LMs.
@@
Table 3: Manticore trained from scratch on LRA using GPT-Neo and Mamba component models. Best accuracies are bolded. GPT-Neo does not support the Pathfinder-X sequence length requirement, so its mixture weight is 0 and Manticore reduces to Mamba.
Setup. Next, we compare Manticore to non-hybrid architectures trained from scratch on the MAD tasks. For these experiments, our component models use the default architecture and training settings used in MAD. We compare two-block GPT-Neo and Mamba models to a Manticore hybrid using a single Manticore block.
Results. Our results are shown in Table 2 (right). Manticore outperforms GPT-Neo and Mamba on all of the MAD tasks in this setting. This provides further evidence for our claim that Manticore hybrids compete with existing LMs when trained from scratch. It is conceivable that our larger Manticore hybrids simply perform better than component models due to their size—however, we find that post-search discretization and retraining tends to result in similar performance, but reduces the model size by roughly half. We include an ablation of post-search discretization in the Appendix.
@@ 4.3 Programming Hybrids
We evaluate C3 with two types of external data: task metadata such as sequence length requirements, and the use of the MAD tasks as a proxy for search on downstream task data.
@@
Setup. As in many of our previous experiments, we used the GPT-Neo and Mamba architectures as component models to our Manticore hybrid. However, this time, we set out to train from scratch on the extremely long-range Pathfinder-X task from LRA, which requires sequence length support greater than that of GPT-Neo. Using this external information about the task, we set the mixture weights for GPT-Neo to 0, which in this case, means that Manticore reduces to Mamba. [^7] Mamba on the LRA is open: https://github.com/state-spaces/mamba/issues/282.
@@
Results. The results of this experiment are shown in the last row of Table 3. In the simple case of having access to task metadata, this validates the claim that we can program mixture weights to exclude incompatible blocks. At the time of writing, we are not aware of prior published Mamba results on LRA despite community interest, which would make our evaluation in Table 3 the first such result. Note that we did not thoroughly tune hyperparameters, so we view this result as a preliminary starting point for the community to build off of, rather than a final answer.
@@
Setup. Finally, in the case in which we can actually run all of our component models on our learning task, we explore when we can program the mixture weights using the MAD tasks as a proxy for search, which are intended to be predictive of scaling laws on The Pile [37, 12]. We set out to fine-tune a pretrained hybrid comprising GPT-Neo-125M and Mamba-130M, which were both pretrained on The Pile, with two Manticore blocks on our Penn Treebank completions synthetic. We train a scaled-down version of this Manticore hybrid with randomly initialized weights and two blocks per component model on the MAD tasks. This yields mixture weights for each of the MAD tasks—we average them across the tasks, and then fine-tune our pretrained hybrid on Penn Treebank completions using the predicted mixture weights.
@@
Results. Our results are shown in Figure 2 (right). We superimpose the predicted mixture weights and mean search trajectory from MAD onto the architecture loss landscape computed on Penn Treebank completions. We find that this procedure recovers a hybrid that outperforms the component models (Mamba, lower right; GPT-Neo, upper left) and substantially outperforms the naive frankenmerges in our search space (upper right and lower left) [13]. This is a scenario in which it is possible to program mixture weights using external sources without performing search on the task data. Intriguingly, search on the MAD tasks appears to follow the architecture gradient on the different downstream fine-tuning task, even though the architecture is scaled-down and trained from scratch on MAD. We hypothesize that programming Manticore hybrids becomes more difficult as the fine-tuning distribution is further from the pretraining distribution, and that the architecture loss landscapes become less similar. This evaluation was carried out on our synthetic PTB completions task, so the fine-tuning dataset should be fairly similar to the pretraining distribution. In our evaluation in Table 1, we find that Mamba outperforms the Pythia component model on English natural language tasks that are further from the pretraining distribution than our synthetic (while both models were trained on The Pile [12] which is largely in English, we are not training on completions produced by the models themselves). Finally, our evaluations in Figure 3 use non-English text, which is further from the pretraining data distribution, and we observe no discernible pattern between their loss landscapes—programming $\alpha$ parameters in this scenario is likely challenging.
5 Conclusions
We present Manticore, a framework that automates the creation of hybrid models from pretrained models by using projectors and convex combinations to align and combine features from multiple different component models, as well as NAS-inspired search procedures. Manticore is efficient and flexible in its usage; hybrids can be trained from scratch, fine-tuned from pretrained component models, and even programmed with external information and/or proxy tasks. In our experiments with several real/synthetic language modeling datasets and existing component/hybrid models, we find that Manticore hybrids match or outperform existing handcrafted hybrid models in these settings. Any requisite fine-tuning and evaluation is performed with a single, large Manticore model rather than designing new hybrids by hand and pretraining them from scratch, which dramatically reduces the computational cost of designing hybrids.
Acknowledgments
We are grateful for the support of the National Science Foundation (NSF) (CCF2106707), the Defense Advanced Research Projects Agency (DARPA Young Faculty Award), and the Wisconsin Alumni Research Foundation (WARF).